14. A compound contains 85.7% by mass of carbon and 14.3%by mass of hydrogen. Its molar mass is 56g/mol. (C= 12, H=1)a) a)Calculate its empirical formula. 


b) Determine its molecular formula.


c) Draw a possible structural formula of the compound.


d) Write a balanced chemical equation for its combustion in oxygen.




15. A sample of lead (II) iodide (an insoluble salt) was prepared by reacting2O cm3 of O.2 mol dm-3 lead (II) nitrate with 40 cm3 of 0.2 mo1 dm-3Potassium iodide


a) What type of reaction occurs between lead (II) nitrate and potassium iodide solutions?


b) Calculate the number of moles of lead (II) nitrate in 20 cm3 of the solution. 


c) Write a balanced equation for the reaction between lead (II) nitrate and potassium iodide solutions.


d) Calculate the mass in grams of lead (II) iodide formed.(Pb=207, I= 127)


1
Expert's answer
2021-02-22T05:55:15-0500

85.7/12 : 14.3/1 = 7,14166667 : 14.3/7,14166667 = 1: 2.

C4H8 = 12 x4 + 1x 8 = 56.

H-C-C-C-C-H

C4H8 + 6O2 = 4CO2 + 4H2O

2. 2KJ + Pb(NO3)2 = PbJ2 +2 KNO3

double-replacement

0.2 x 20/1000 = 0,004 moles.

0.004 x 461 = 1,844.